A pulley weighing 1000 N is supported on a horizontal shaft of negligible weight between two bearings, A and B, 5 metres apart. The pulley is fixed at a distance of 2 m from the bearing A. The mass centre of the polley is 20 mm from the axis of the shaft. Determine the maximum pressure on the two bearings when the shaft rotates at 200 r.p.m.
Neglect the effect of deflection of the axis of the shaft.
1
Expert's answer
2021-01-05T12:05:03-0500

"F=mv^2/r\to F=(mg)\omega^2r/g=1000\cdot(200/60)^2\cdot0.02/9.8=1360(N)" . Total force.


"5R-1360\cdot0.02=0\to R=514(N)" . Force on bearing B


"1360-514=816(N)" . Force on bearing A